\begin{references}

\bibitem{SK}
B. Z. Spivak and S. A. Kivelson,
Phys.\ Rev. B {\bf 43}, 3740 (1991).

\bibitem{Bulaevskii}
L. N. Bulaevskii, V. V. Kuzii, and A. A. Sobyanin,
JETP\ Lett. {\bf 25}, 290 (1977).

\bibitem{KS} 
S. A. Kivelson and B. Z. Spivak,
Phys.\ Rev.\ B {\bf 45}, 10490 (1992).

\bibitem{Glazman} 
L. I. Glazman and K. A. Matveev, 
JETP\ Lett. {\bf 49}, 659 (1989).

\bibitem{footnote1} 
The sign of the supercurrent can be switched quite easily
by bringing the Josephson junction out of equilibrium,
see J. J. A. Baselmans, A. F. Morpurgo,
B. J. van Wees, T. M. Klapwijk, Nature {\bf 397}, 43 (1999).
This is not a $\pi$-junction in the original sense,
which refers to an equilibrium property of the junction.

\bibitem{footnote2} 
The sign of the supercurrent becomes entirely sample specific
if time-reversal symmetry is broken by a magnetic field or magnetic
order, see F. Zhou, Int.\ J.\ Mod.\ Phys.\ B
{\bf 13}, 2229 (1999).

\bibitem{Altshuler}
B. L. Altshuler and B. Z. Spivak,
Sov.\ Phys.\ JETP {\bf 65}, 343 (1987).

\bibitem{Bezryadin}
A. Bezryadin, C. N. Lan, and M. Tinkham,
Nature {\bf 404}, 971 (2000).

\bibitem{BB}
P. W. Brouwer and C. W. J. Beenakker,
Chaos, Solitons \& Fractals
{\bf 8}, 1249 (1997).
A factor of $2$ is missing in Eq.\ (8)
of this publication. The preprint 
cond-mat/9611162  does not contain this misprint.


\end{references}

\documentstyle[aps,prb,twocolumn,epsfig,floats,amssymb,amsfonts]{revtex}

\def\s{\sum\limits}

\def\be{\begin{equation}}
\def\e{\end{equation}}
\def\beml{\begin{mathletters}}
\def\eml{\end{mathletters}}
\def\beq{\begin{eqnarray}}
\def\eq{\end{eqnarray}}
\def\ba{\begin{array}}
\def\a{\end{array}}
\def\d{\dagger}
\def\l{\left}
\def\r{\right}
\def\n{\nonumber}
\def\la{\langle}
\def\ra{\rangle}
\def\det{\,{\rm Det}\,}
\def\tr{{\rm Tr}\,}

\def\im{\,{\rm Im}\,}
\def\re{\,{\rm Re}\,}

\def\ep{\varepsilon}
\begin{document}
 
\draft

\date{May 2000}

\title{Can mesoscopic fluctuations
reverse the supercurrent through a disordered 
Josephson junction?}

\author{M. Titov, Ph. Jacquod, and C. W. J. Beenakker}

\address{
Instituut-Lorentz, Universiteit Leiden,
P.\,O.~Box 9506, 2300 RA Leiden,
The Netherlands
}
\twocolumn[
\widetext
\begin{@twocolumnfalse}

\maketitle

\begin{abstract}
We calculate the Josephson coupling energy 
$U_J(\phi)$ (related to the supercurrent
$I=(2e/\hbar) dU_J/d\phi$) for a disordered normal metal
between two superconductors with a phase difference $\phi$.
We demonstrate that the symmetry of the scattering matrix
of non-interacting quasiparticles in zero magnetic field
implies that $U_J(\phi)$ has a minimum at $\phi=0$.
A maximum (that would lead to a $\pi$-junction or negative
superfluid density) is excluded for any realization
of the disorder.
\end{abstract}
\pacs{PACS numbers: 74.80.Fp, 72.15.Rn, 73.63.Rt, 74.40.+k}

\end{@twocolumnfalse}
]
\narrowtext

The question in the title was posed ten years ago by Spivak
and Kivelson,\cite{SK} in their search for
mechanisms that would lead to a negative local superfluid density
in a dirty superconductor. A negative instead of a positive 
superfluid density means that the Josephson coupling energy 
$U_J(\phi)$ has a maximum instead of a minimum for zero
phase difference $\phi$ of the superconducting order parameter.
The supercurrent $I(\phi)=(2e/\hbar) dU_J/d\phi$ then has the opposite sign
as usual for small $\phi$. This has a variety of observable consequences,
including a ground state with a non-zero supercurrent,\cite{Bulaevskii}
Aharonov-Bohm oscillations with period $h/4 e$, and negative
magnetoresistance.\cite{KS} A Josephson junction with a negative superfluid density
is known as a $\pi$-junction,\cite{Bulaevskii} because the ground state for
large magnetic inductance is close to $\phi=\pi$ instead of 
as usual at $\phi=0$.

The known mechanism for the creation of a $\pi$-junction\cite{SK,Bulaevskii,KS,Glazman}
in equilibrium\cite{footnote1} in a non-magnetic material\cite{footnote2}
requires strong Coulomb repulsion to create a localized spin on a resonant impurity 
level. Spivak and Kivelson asked the question whether purely 
one-electron conductance fluctuations might be sufficient to produce a locally negative
superfluid density near the insulating state. A suggestive argument that
this might be possible comes from the relative magnitude
of the mesoscopic sample-to-sample fluctuations of the supercurrent in
a disordered superconductor--normal-metal--superconductor
(SNS) junction.\cite{Altshuler}
The ratio $\la \delta I^2 \ra^{1/2}/\la I\ra \simeq e^2/h G$ of the 
root-mean-squared fluctuations over the mean supercurrent is 
$\ll 1$ if the conductance $G$ of the normal metal is large
compared to the conductance quantum $e^2/h$. The ratio becomes 
of order unity on approaching the insulating state, suggesting
that the supercurrent might have a negative value in some samples.

Of course, the root-mean-square amplitude of the supercurrent fluctuations
does not
distinguish between a positive and negative sign, so that this argument 
is only suggestive. We were motivated to settle this issue
because of recent experiments on localization in quasi-one-dimensional 
superconductors.\cite{Bezryadin}
This has renewed the interest in the fundamental question whether mesoscopic 
fluctuations are sufficient or not to create a negative local superfluid
density. The answer, as we will show, is that they are not.

The two superconductors that form the SNS junction
have
order parameters $\Delta e^{i\phi/2}$ and $\Delta e^{-i\phi/2}$.
The contacts to the normal metal 
have $N$ propagating modes at the Fermi energy $E_F$, so that the elastic
scattering
by the normal metal at energy $E=E_F+\ep$ is characterized by a 
$2N\times 2N$
scattering matrix $S(\ep)$. The two properties of $S$ that we use
are that it is analytic in the upper half of the complex $\ep$ plane
and that it is a symmetric matrix, $S(\ep)=S(\ep)^T$,
when time-reversal symmetry is preserved. 

The starting point of our calculation is the relationship 
derived in Ref.\ \onlinecite{BB}
between 
the Josephson coupling energy  $U_J(\phi)$ in equilibrium 
at temperature $T$ and the scattering matrix,
\be
\label{basic}
U_J= -2 k_B T \s_{n=0}^\infty \ln \det 
\l[1-S_A(i\omega_n) S_N(i\omega_n) \r].
\e
The summation runs over the Matsubara frequencies 
$\omega_n= (2 n+1)\pi k_B T$.
The $4N\times 4N$ scattering matrix $S_N(\ep)$ describes
the elastic scattering by disorder in the normal metal
of non-interacting electron and hole quasiparticles
with excitation energy $\ep$,
\be
\label{SN}
S_N(\ep)=\l(
\ba{cc} S(\ep) & 0\\ 0 & S(-\ep)^* \a
\r).
\e 
The analytical continuation of $S_{N}$ from real 
to imaginary energy ($\varepsilon\rightarrow i\omega$) 
follows from $S(\varepsilon)\rightarrow S(i\omega)$ and
$S(-\varepsilon)^{\ast}\rightarrow S(i\omega)^{\ast}$.
Similarly, the matrix $S_A(\ep)$ describes the 
Andreev reflection from the superconductors,
\beml
\label{SA}
\beq
\label{SAa}
S_A(\ep)&=&\alpha(\ep)
\l( \ba{cc} 0 & e^{i \Lambda \phi/2} \\
e^{-i \Lambda \phi/2} & 0
\a \r),\\ 
\label{SAb}
\alpha(\ep)&=&\ep/\Delta-
i\sqrt{1-\ep^2/\Delta^2}.
\eq
\eml
Here $\Lambda$ is a $2N\times 2N$ diagonal matrix
with elements $\Lambda_{jj}=1$ for
$1\le j \le N$ and $\Lambda_{jj}=-1$ for $N+1\le j\le 2N$.

Eq.\ (\ref{basic}) differs from the usual 
representation of the Josephson energy as a sum over the
discrete spectrum $(\ep<\Delta)$ plus an integration over
the continuous spectrum $(\ep > \Delta)$. 
Each term in the sum (\ref{basic}) is combined from
contributions out of the discrete and the continuous spectrum.
We will now show that each of these combinations is minimal
for $\phi=0$, although the contributions from 
the discrete and continuous spectrum separately are not.

Let us abbreviate
\be
Z(\omega)=(\sqrt{1+\omega^2/\Delta^2}
-\omega/\Delta)e^{-i \Lambda \phi/2} S(i\omega).
\e
Using the identity $\ln \det =\tr \ln$ in Eq.\ (\ref{basic}) 
one can calculate the first and second derivative 
with respect to $\phi$ of the energy $U_J(\phi)$. 
The first derivative is given by
\be
\label{firsta}
{d U_J\over d\phi} = 2 k_B T \s_{n=0}^\infty \im \tr 
\l[ h_{11}(\omega_n) -h_{22}(\omega_n) \r],
\e
where the $N\times N$ matrices $h_{11}$ and $h_{22}$
are blocks of the matrix
\be
\label{firstb}
H=Z^* Z(1+Z^* Z)^{-1}=\l(
\ba{cc} h_{11}& h_{12}\\ h_{21} & h_{22} \a
\r).
\e
The first derivative
is equal to the supercurrent and vanishes at $\phi=0$ 
as dictated by time-reversal symmetry.

For the second derivative we obtain 
\beq
{d^2 U_J\over d\phi^2}& =& 4 k_B T \s_{n=0}^\infty \re \tr 
\l[ f_{12}(\omega_n) f_{21}^*(\omega_n) \r.\n \\
\label{seconda}
&& \l. +h_{12}(\omega_n) h_{21}(\omega_n) \r],\\
\label{secondb}
F&=&Z(1+Z^* Z)^{-1}=\l(
\ba{cc} f_{11}& f_{12}\\ f_{21} & f_{22} \a
\r).
\eq
At $\phi=0$ the symmetry of $S$ implies that 
$F=F^T$ and $H=H^\d$, hence  $f_{21}^*=f_{12}^\d$, $h_{21}=h_{12}^\d$.
Therefore, every term in the sum (\ref{seconda})
is positive. We conclude that the
Josephson energy
$U_J(\phi)$ has a minimum at $\phi=0$:
\be
\l.{dU_J\over d\phi}\r|_{\phi=0}=0, \qquad
\l.{d^2U_J\over d\phi^2}\r|_{\phi=0}>0.
\e

This concludes the proof that mesoscopic fluctuations can not invert the
stability of the SNS junction at zero phase, excluding a mechanism for
the creation of a $\pi$-junction proposed ten years ago.\cite{SK} 
The proof holds for non-interacting
quasiparticles in zero magnetic field at arbitrary temperature, for any
disorder potential and any dimensionality of the junction. As a final
remark, we conjecture (and have a proof for $N=1$) that Eq.\ (1) implies
$dU_{J}/d\phi>0$ in the entire interval $0<\phi<\pi$ in the presence of
time-reversal symmetry.

We thank Boris Spivak for urging us to solve this problem and for valuable
discussions. This work was supported by the Dutch Science Foundation NWO/FOM.

\end{document}